# NPHardEval: Dynamic Benchmark on Reasoning Ability of Large Language Models via Complexity Classes

### 22 Dec 2023 v1

@@ 5.3 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
Fewshot (-5) 0.977 0.685 0.920 0.735 0.603 0.000 0.025 0.000 0.000 0.122
Fewshot (-4) 1.000 0.662 0.902 0.667 0.516 0.053 0.060 0.000 0.000 0.073
Fewshot (-3) 0.982 0.694 0.831 0.769 0.496 0.029 0.084 0.031 0.000 0.129
Fewshot (-2) 1.000 0.771 0.910 0.710 0.617 0.054 0.094 0.044 0.000 0.063
Fewshot (-1) 0.987 0.770 0.896 0.589 0.607 0.048 0.124 0.043 0.004 0.069
Fewshot (0) 0.984 0.671 0.845 0.651 0.660 0.104 0.220 0.065 0.009 0.124
Fewshot (1) 0.991 0.580 0.878 0.696 0.455 0.138 0.140 0.076 0.000 0.089
Fewshot (2) 1.000 0.675 0.829 0.587 0.656 0.216 0.164 0.071 0.015 0.040
Fewshot (3) 0.993 0.736 0.800 0.598 0.489 0.173 0.187 0.018 0.035 0.109
Fewshot (4) 1.000 0.729 0.869 0.580 0.471 0.176 0.135 0.022 0.035 0.082
Fewshot (5) 1.000 0.671 0.844 0.602 0.607 0.167 0.227 0.027 0.007 0.100
@@
Fewshot (-5) 0.000 0.060 0.020 0.000 0.000 1.000 0.980 1.000 1.000 0.900
Fewshot (-4) 0.000 0.033 0.000 0.000 0.017 0.950 0.933 1.000 1.000 0.933
Fewshot (-3) 0.000 0.057 0.057 0.014 0.000 0.957 0.914 0.971 1.000 0.857
Fewshot (-2) 0.000 0.075 0.025 0.000 0.013 0.925 0.888 0.950 0.975 0.913
Fewshot (-1) 0.000 0.044 0.022 0.000 0.011 0.911 0.867 0.922 0.922 0.933
Fewshot (0) 0.000 0.060 0.020 0.000 0.020 0.800 0.760 0.810 0.930 0.850
Fewshot (1) 0.000 0.040 0.020 0.010 0.010 0.760 0.790 0.850 1.000 0.880
Fewshot (2) 0.000 0.040 0.040 0.000 0.010 0.650 0.830 0.830 0.980 0.900
Fewshot (3) 0.000 0.020 0.050 0.010 0.030 0.680 0.770 0.890 0.930 0.880
Fewshot (4) 0.000 0.050 0.030 0.000 0.040 0.660 0.800 0.900 0.930 0.870
Fewshot (5) 0.000 0.050 0.040 0.000 0.020 0.680 0.720 0.950 0.960 0.890

### 25 Dec 2023 v2

@@ 5.3 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
Fewshot (-5) 0.978 0.685 0.920 0.735 0.603 0.065 0.043 0.165 0.085 0.155
Fewshot (-4) 1.000 0.662 0.902 0.667 0.516 0.093 0.129 0.156 0.084 0.118
Fewshot (-3) 0.982 0.694 0.831 0.769 0.496 0.143 0.153 0.131 0.067 0.084
Fewshot (-2) 1.000 0.771 0.910 0.710 0.617 0.102 0.094 0.117 0.037 0.087
Fewshot (-1) 0.987 0.770 0.896 0.589 0.607 0.048 0.085 0.157 0.087 0.057
Fewshot (0) 0.984 0.671 0.846 0.651 0.660 0.598 0.413 0.258 0.089 0.098
Fewshot (1) 0.991 0.580 0.878 0.696 0.455 0.593 0.386 0.233 0.055 0.109
Fewshot (2) 1.000 0.675 0.829 0.587 0.656 0.647 0.296 0.175 0.067 0.056
Fewshot (3) 0.993 0.736 0.800 0.598 0.489 0.662 0.318 0.144 0.093 0.098
Fewshot (4) 1.000 0.729 0.869 0.580 0.471 0.638 0.287 0.269 0.053 0.106
Fewshot (5) 1.000 0.671 0.844 0.602 0.607 0.167 0.356 0.202 0.055 0.064
@@
Fewshot (-5) 0.000 0.060 0.020 0.000 0.000 0.940 0.480 0.160 0.640 0.860
Fewshot (-4) 0.000 0.033 0.000 0.000 0.017 0.917 0.617 0.117 0.633 0.900
Fewshot (-3) 0.000 0.057 0.057 0.014 0.000 0.871 0.471 0.043 0.600 0.914
Fewshot (-2) 0.000 0.075 0.025 0.000 0.013 0.888 0.538 0.063 0.613 0.888
Fewshot (-1) 0.000 0.044 0.022 0.000 0.011 0.911 0.622 0.078 0.589 0.933
Fewshot (0) 0.000 0.060 0.020 0.000 0.020 0.300 0.380 0.040 0.540 0.880
Fewshot (1) 0.000 0.040 0.020 0.010 0.010 0.290 0.420 0.060 0.520 0.830
Fewshot (2) 0.000 0.040 0.040 0.000 0.010 0.290 0.440 0.030 0.640 0.910
Fewshot (3) 0.000 0.020 0.050 0.010 0.030 0.280 0.460 0.030 0.650 0.830
Fewshot (4) 0.000 0.050 0.030 0.000 0.040 0.280 0.530 0.080 0.630 0.850
Fewshot (5) 0.000 0.050 0.040 0.000 0.020 0.680 0.440 0.050 0.650 0.900